Project

General

Profile

2019-11-01 ADR Cryostat Run 03 Notes

Overview

This run was intended as the third test of the optical filtering configuration of the ADR cryostat. In the second test, the filtering stack was as follows (in order of decreasing :

  • 300K:
    • zotefoam window
  • 50K:
    • 6x? 1/8" thick zotefoam shaders
    • 8 icm Cardiff filter with mupor AR coating
  • 4K:
    • 9 icm Cardiff filter

If this configuration seems weird, it's because it is. The 8 icm filter is absorbing basically the entire 300K blackbody spectrum, so probably heating up well above 50K, then re-radiating onto a 9 icm filter that is absorbing its entire spectrum. The purpose of this run was to test a new filtering stack that should result in significantly lower temperatures at 4K.

New filter stack

The new filtering stack that I installed consisted of the following:

  • 300K:
    • zotefoam window
  • 50K:
    • 6x? 1/8" thick zotefoam shaders
    • 3/16" HDPE
  • 4K:
    • 3/16" nylon (on inside of 4K volume)
    • 9 icm Cardiff filter (on inside of 4K volume)

Some pictures show the filtering configuration below:


Figure 1: 4K plate, showing nylon filter and spacer ring that separates the nylon filter from the 9 icm Cardiff filter.


Figure 2: Side view of 4K plate, showing nylon filter (thickest), spacer ring, and 9 icm Cardiff filter, after clamping ring was installed, but before taping over with Al.


Figure 3: Finalized 4K plate just before installation, showing position of diode thermometer.


Figure 4: Finalized 50K plate just before installation, showing HDPE filter

2019-10-04 thermal tests

We reached base temperature over the weekend. On Monday morning at 9:45am, the temperatures were:
  • 4K LPF: 7.827K
  • 3K PTC stage: 3.1K
  • 50K PTC stage: 59.1K

I ran an ADR cycle and the hold time was about 150 minutes.

In the evening, I ran two more fridge cycles, which basically did not stay cold, although it is unclear whether this was due to a thermal issue or whether the PID regulation was not engaging correctly.

2019-10-05 thermal tests

When I arrived in the morning, the 4K filter was 7.860K, similar to yesterday. I initially turned off the PTC for several minutes, but had a change of heart and decided to run one more fridge cycle to see whether the two cycles of last night with no hold time were compromised by a problem with the PID not engaging early enough. In addition, I wanted to take some time to digest the data from the cycles (see attached adr-log-2019-10-28.csv).

Figure 5 below shows the FAA temperature and magnet current during the first two cycles run yesterday. The first cycle achieved a hold time of about 150min. The second cycle didn't really have much of a hold time, although as Figure 5 indicates, this is largely due to the fact that the PID didn't engage until after nearly 1hr. Note that in both cases, the minimum temperature achieved by the FAA was about 135mK. This suggests that the thermal performance was about the same in both cases.